# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 3.2.1   Spin-flavor $SU(6)$
Table 3.1: $SU(3)_{f}$ flavor wave functions for octet, decuplet and singlet baryons. $I$ and $I_{3}$ denote the isospin and isospin-3 component and S the strangeness.
  Flavor content Baryon $I$ $I_{3}$ S Flavor wave function
  $uud$ $p$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 00 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}udu-duu\\ -\frac{1}{\sqrt{3}}\,(udu+duu-2uud)\end{array}\right]$
$udd$ $n$ $\nicefrac{{1}}{{2}}$ $-\nicefrac{{1}}{{2}}$ 00 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}udd-dud\\ \frac{1}{\sqrt{3}}\,(dud+udd-2ddu)\end{array}\right]$
  $uus$ $\Sigma^{+}$ $1$ $1$ $-1$ $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}usu-suu\\ -\frac{1}{\sqrt{3}}\,(usu+suu-2uus)\end{array}\right]$
$uds$ $\Sigma^{0}$ $1$ 00 $-1$ $\frac{1}{2}\left[\begin{array}[]{c}sud-usd+sdu-dsu\\ \frac{1}{\sqrt{3}}\,(sud+usd+sdu+dsu-2uds-2dus)\end{array}\right]$
$dds$ $\Sigma^{-}$ $1$ $-1$ $-1$ $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}dsd-sdd\\ -\frac{1}{\sqrt{3}}\,(dsd+sdd-2dds)\end{array}\right]$
  $uds$ $\Lambda^{0}$ 00 00 $-1$ $\frac{1}{2}\left[\begin{array}[]{c}\frac{1}{\sqrt{3}}\,(2uds-2dus+usd-dsu+sdu-sud)\\ usd-dsu+sud-sdu\end{array}\right]$
  $uss$ $\Xi^{0}$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ $-2$ $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}uss-sus\\ \frac{1}{\sqrt{3}}\,(sus+uss-2ssu)\end{array}\right]$
$dss$ $\Xi^{-}$ $\nicefrac{{1}}{{2}}$ $-\nicefrac{{1}}{{2}}$ $-2$ $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}dss-sds\\ \frac{1}{\sqrt{3}}\,(sds+dss-2ssd)\end{array}\right]$
  $uuu$ $\Delta^{++}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{3}}{{2}}$ 00 $uuu$
$uud$ $\Delta^{+}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{1}}{{2}}$ 00 $\frac{1}{\sqrt{3}}\,(uud+udu+duu)$
$udd$ $\Delta^{0}$ $\nicefrac{{3}}{{2}}$ $-\nicefrac{{1}}{{2}}$ 00 $\frac{1}{\sqrt{3}}\,(udd+dud+ddu)$
$ddd$ $\Delta^{-}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{3}}{{2}}$ 00 $ddd$
  $uus$ $\Sigma^{+}$ $1$ $1$ $-1$ $\frac{1}{\sqrt{3}}\,(uus+usu+suu)$
$uds$ $\Sigma^{0}$ $1$ 00 $-1$ $\frac{1}{\sqrt{6}}\,(uds+sud+dsu+dus+usd+sdu)$
$dds$ $\Sigma^{-}$ $1$ $-1$ $-1$ $\frac{1}{\sqrt{3}}\,(dds+dsd+sdd)$
  $uss$ $\Xi^{0}$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ $-2$ $\frac{1}{\sqrt{3}}\,(uss+sus+ssu)$
$dss$ $\Xi^{-}$ $\nicefrac{{1}}{{2}}$ $-\nicefrac{{1}}{{2}}$ $-2$ $\frac{1}{\sqrt{3}}\,(dss+sds+ssd)$
  $sss$ $\Omega^{-}$ 00 00 $-3$ $sss$
  $uds$ $\Lambda^{0}$ 00 00 $-1$ $\frac{1}{\sqrt{6}}\,(uds+sud+dsu-dus-usd-sdu)$
In the non-relativistic quark model one goes a step further by combining the $SU(2)_{s}$ spin and $SU(3)_{f}$ flavor wave functions into $SU(2)_{s}\otimes SU(3)_{f}\simeq SU(6)$ representations. The spins $\bm{s}_{i}$ of the three quarks combine to form the total quark spin $\bm{S}$ with $S=\nicefrac{{1}}{{2}}$ or $\nicefrac{{3}}{{2}}$. The $SU(2)_{s}$ spin wave functions, with index $s$ for spin, can be read off from the entries with $uuu$, $uud$, $udd$, $ddd$ in Table 3.1 with the replacements $u\to$ spin up ($\uparrow$) and $d\to$ spin down ($\downarrow$). This yields two doublets for total quark spin $S=\frac{1}{2}$,
@@
$\begin{split}\mathcal{S}_{s}\,=\;&|\uparrow\uparrow\uparrow\,\rangle\,,\quad\frac{1}{\sqrt{3}}\,|\uparrow\uparrow\downarrow+\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow\,\rangle\,,\quad\frac{1}{\sqrt{3}}\,|\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow+\downarrow\downarrow\uparrow\,\rangle\,,\quad|\downarrow\downarrow\downarrow\,\rangle\,.\end{split}$ (3.6)
@@
It is instructive to relate these wave functions with the underlying diquark structure. A diquark wave function satisfies similar constraints as the baryon wave function: it must be antisymmetric under quark exchange due to the Pauli principle, and because the color part is antisymmetric, the remainder must be symmetric. Combining $u$ and $d$ quarks yields an antisymmetric isospin singlet $\frac{1}{\sqrt{2}}(ud-du)$ with $I=0$ and a symmetric triplet $\{uu,\frac{1}{\sqrt{2}}(ud+du),dd\}$ with $I=1$. These must be combined with spin parts of the same symmetry. For ground states with $L=0$, one arrives at scalar diquarks ($J^{P}=0^{+}$) with $S=0$, $I=0$ and axialvector diquarks ($J^{P}=1^{+}$) with $S=1$, $I=1$. These were nicknamed “good” and “bad” diquarks by Wilczek [519] and they play a significant role; e.g., in diquark models axialvector diquarks are typically 200 …300 MeV heavier than scalar diquarks [520, 521, 522]. Because of isospin, the ground-state nucleon ($I=0$) contains both scalar and axialvector diquarks whereas the $\Delta$ resonance ($I=1$) consists only of axialvector diquarks. The diquark components enter in the spin and flavor wave functions of baryons (Fig. 3.1): the upper ($\mathcal{M}_{\mathcal{A}}$) components in the doublets are antisymmetric in the first two indices and can thus be related to the “good” diquarks, whereas the lower ($\mathcal{M}_{\mathcal{S}}$) components and the singlets $\mathcal{S}$ are symmetric and associated with the “bad” diquarks. This will play a role in the quark models discussed below.
@@ 3.2.2    Spatial wave functions and the non-relativistic harmonic oscillator
To classify the baryon spectrum, we must also discuss the spatial or orbital wave functions $|\psi_{\rm space}\rangle$. The three-body problem is characterized by three degrees of freedom. Instead of the position vectors $\bm{x}_{i}\,(i=1,2,3)$ of the three particles, one usually defines the Jacobi coordinates
@@
which include the trivial center-of-mass (CM) motion and two oscillations represented by the variables $\bm{\rho}$ and $\bm{\lambda}$. The $\bm{\rho}$ oscillator describes the dynamics within a diquark and the $\bm{\lambda}$ oscillator the motion of the third quark relative to the diquark. Permutation symmetry ensures that all three quarks have identical functions. Instead of Jacobian coordinates, some authors [523, 524, 525, 526, 527, 528, 529, 530, 531, 532, 533] use hypercentral coordinates given by the collective size variable $\sqrt{\rho^{2}+\lambda^{2}}$ and an angle $\arctan\frac{\rho}{\lambda}$.
In a series of papers, Isgur and Karl studied masses and mixing angles of low-mass baryons in a model that contains a flavor-independent confinement potential and hyperfine interactions between quarks due to one-gluon exchange [534, 535, 536, 537, 538, 539, 540]. The confining potential is assumed to be harmonic, and the deviation due to the linearly rising confining potential expected in QCD is treated perturbatively. In our following interpretation of the baryon resonance spectrum, we will often use the symmetry properties of the nonrelativistic quark model with a harmonic oscillator as confining potential since it predicts the number of states and, at least qualitatively, their masses. The eigenfunctions of the Hamiltonian are then expanded into eigenfunctions of the harmonic oscillator serving as a complete orthonormal set of basis functions.
@@
Figure 3.2: (color online) Light baryon spectrum in the harmonic oscillator model with band quantum number N, spin $S$, orbital angular momentum $L$, total angular momentum $J$ and parity $P$. $F$ denotes the flavor multiplet ($N$ for octet, $\Delta$ for decuplet) and $\mathbf{D}$ the spin-flavor multiplet. The assignment with experimental states is discussed in the text. Three- and four-star resonances are shown in bold font. the $J^{P}$ quantum numbers in gray do not have good experimental candidates. The background colors refer to the orbital angular momentum $L$.
The wave functions within a given shell can be further rearranged to form combinations with definite S3 symmetry, as shown in the right panel of Fig. 3.1 (see e.g. Ref. [541] for details). This yields the spatial wave functions $\mathcal{S}_{o}$, $\mathcal{D}_{o}$ and $\mathcal{A}_{o}$ appearing in the left panel of Fig. 3.1. The total orbital angular momentum $\bm{L}=\bm{l}_{\rho}+\bm{l}_{\lambda}$ can adopt values $L=|l_{\rho}-l_{\lambda}|,\dots,l_{\rho}+l_{\lambda}$, and $\bm{L}$ combines with the quark spin $\bm{S}$ to the spin of the particle $\bm{J}=\bm{L}+\bm{S}$. The quantum numbers $J^{P}$ are then obtained from the parity $P=(-1)^{l}$.
@@ 3.3   Light baryon spectrum
Figure 3.4: The number of 3* and 4* $N^{*}$ and $\Delta^{*}$ resonances in [3] above the ground states in 50 MeV bins. The arrows mark the largest spin up to the given mass.
In each cluster there is a state with maximal spin. These states are referred to as stretched states, where the total angular momentum is $J=L+S$. In the high-mass region, the stretched states dominate the reaction, either on physical grounds or because the largest partial wave is the easiest one to identify in angular distributions. The left panels in Fig. 3.5 show the squared baryon mass with maximal spin $J$ in a given cluster, for positive-parity $N^{*}$ and $\Delta^{*}$ resonances, respectively, as a function of $J$. One observes a Regge behavior, i.e., the squared masses of the stretched states depend linearly on $J$ [542]. Similar Regge trajectories can be drawn for isoscalar mesons, including the $\omega$, $f_{2}(1270)$, $\omega_{3}(1670)$, $f_{4}(1950)$, $\rho_{5}(2350)$, and $f_{6}(2510)$. Mesons and $\Delta^{*}$ baryons have very similar Regge slopes: writing $M^{2}=M_{0}^{2}+aJ$, one finds $a=1.14(2)$ GeV2 for mesons and $a=0.98(2)$ GeV2 for baryons. In mesons, the quark and antiquark are bound by the strong force between a colored quark and an antiquark carrying anti-color. Two colored quarks in a baryon can form a diquark carrying the anti-color of the third quark. The similarity of the meson and baryon trajectories has led to the hypothesis that baryon resonances (more precisely, the stretched states) can be understood as quark-diquark oscillations.
In the following we assign the $N^{*}$ and $\Delta^{*}$ resonances to $SU(6)$ multiplets with defined shell number N, intrinsic orbital angular momentum $L$, total quark spin $S$ and parity $P$. To this end, we combine the masses of baryons in the same multiplet (which differ only by $J$) to the center-of-gravity (c.o.g.) mass [^1] The c.o.g. mass can be used for heavy quarkonia to estimate the mass of the singlet meson from the three triplet states, provided spin-spin and tensor interactions are small. It has been applied to light-quark mesons [543] even though this is not really justified. The idea is to find a quantity that does not dependent on spin-orbit interactions. defined by

### v4

@@ 3.2.1  Spin-flavor $SU(6)$
Table 3.1: $SU(3)_{f}$ flavor wave functions for octet, decuplet and singlet baryons. II and $I_{3}$ denote the isospin and isospin-3 component and S the strangeness.
  Flavor content Baryon II $I_{3}$ S Flavor wave function
  uuduud $p$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 0 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}udu-duu\\ -\frac{1}{\sqrt{3}}\,(udu+duu-2uud)\end{array}\right]$
uddudd $n$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 0 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}udd-dud\\ \frac{1}{\sqrt{3}}\,(dud+udd-2ddu)\end{array}\right]$
  uusuus $\Sigma^{+}$ 11 11 1-1 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}usu-suu\\ -\frac{1}{\sqrt{3}}\,(usu+suu-2uus)\end{array}\right]$
udsuds $\Sigma^{0}$ 11 0 1-1 $\frac{1}{2}\left[\begin{array}[]{c}sud-usd+sdu-dsu\\ \frac{1}{\sqrt{3}}\,(sud+usd+sdu+dsu-2uds-2dus)\end{array}\right]$
ddsdds $\Sigma^{-}$ 11 1-1 1-1 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}dsd-sdd\\ -\frac{1}{\sqrt{3}}\,(dsd+sdd-2dds)\end{array}\right]$
  udsuds $\Lambda^{0}$ 0 0 1-1 $\frac{1}{2}\left[\begin{array}[]{c}\frac{1}{\sqrt{3}}\,(2uds-2dus+usd-dsu+sdu-sud)\\ usd-dsu+sud-sdu\end{array}\right]$
  ussuss $\Xi^{0}$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 2-2 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}uss-sus\\ \frac{1}{\sqrt{3}}\,(sus+uss-2ssu)\end{array}\right]$
dssdss $\Xi^{-}$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 2-2 $\frac{1}{\sqrt{2}}\left[\begin{array}[]{c}dss-sds\\ \frac{1}{\sqrt{3}}\,(sds+dss-2ssd)\end{array}\right]$
  uuuuuu $\Delta^{++}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{3}}{{2}}$ 0 uuuuuu
uuduud $\Delta^{+}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{1}}{{2}}$ 0 $\frac{1}{\sqrt{3}}\,(uud+udu+duu)$
uddudd $\Delta^{0}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{1}}{{2}}$ 0 $\frac{1}{\sqrt{3}}\,(udd+dud+ddu)$
dddddd $\Delta^{-}$ $\nicefrac{{3}}{{2}}$ $\nicefrac{{3}}{{2}}$ 0 dddddd
  uusuus $\Sigma^{+}$ 11 11 1-1 $\frac{1}{\sqrt{3}}\,(uus+usu+suu)$
udsuds $\Sigma^{0}$ 11 0 1-1 $\frac{1}{\sqrt{6}}\,(uds+sud+dsu+dus+usd+sdu)$
ddsdds $\Sigma^{-}$ 11 1-1 1-1 $\frac{1}{\sqrt{3}}\,(dds+dsd+sdd)$
  ussuss $\Xi^{0}$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 2-2 $\frac{1}{\sqrt{3}}\,(uss+sus+ssu)$
dssdss $\Xi^{-}$ $\nicefrac{{1}}{{2}}$ $\nicefrac{{1}}{{2}}$ 2-2 $\frac{1}{\sqrt{3}}\,(dss+sds+ssd)$
  ssssss $\Omega^{-}$ 0 0 3-3 ssssss
  udsuds $\Lambda^{0}$ 0 0 1-1 $\frac{1}{\sqrt{6}}\,(uds+sud+dsu-dus-usd-sdu)$
In the non-relativistic quark model one goes a step further by combining the $SU(2)_{s}$ spin and $SU(3)_{f}$ flavor wave functions into $SU(2)_{s}\otimes SU(3)_{f}\simeq SU(6)$ representations. The spins $\bm{s}_{i}$ of the three quarks combine to form the total quark spin $\bm{S}$ with $S=\nicefrac{{1}}{{2}}$ or $\nicefrac{{3}}{{2}}$. The $SU(2)_{s}$ spin wave functions, with index $s$ for spin, can be read off from the entries with uuuuuu, uuduud, uddudd, dddddd in Table 3.1 with the replacements $u\to$ spin up ($\uparrow$) and $d\to$ spin down ($\downarrow$). This yields two doublets for total quark spin $S=\frac{1}{2}$,
@@
[…]
@@
It is instructive to relate these wave functions with the underlying diquark structure. A diquark wave function satisfies similar constraints as the baryon wave function: it must be antisymmetric under quark exchange due to the Pauli principle, and because the color part is antisymmetric, the remainder must be symmetric. Combining $u$ and $d$ quarks yields an antisymmetric isospin singlet $\frac{1}{\sqrt{2}}(ud-du)$ with $I=0$ and a symmetric triplet $\{uu,\frac{1}{\sqrt{2}}(ud+du),dd\}$ with $I=1$. These must be combined with spin parts of the same symmetry. For ground states with $L=0$, one arrives at scalar diquarks ($J^{P}=0^{+}$) with $S=0$, $I=0$ and axialvector diquarks ($J^{P}=1^{+}$) with $S=1$, $I=1$. These were nicknamed “good” and “bad” diquarks by Wilczek Wilczek:2004im and they play a significant role; e.g., in diquark models axialvector diquarks are typically 200 …300 MeV heavier than scalar diquarks Anselmino:1992vg , Santopinto:2014opa , Barabanov:2020jvn . Because of isospin, the ground-state nucleon ($I=0$) contains both scalar and axialvector diquarks whereas the $\Delta$ resonance ($I=1$) consists only of axialvector diquarks. The diquark components enter in the spin and flavor wave functions of baryons (Fig. 3.1): the upper ($\mathcal{M}_{\mathcal{A}}$) components in the doublets are antisymmetric in the first two indices and can thus be related to the “good” diquarks, whereas the lower ($\mathcal{M}_{\mathcal{S}}$) components and the singlets $\mathcal{S}$ are symmetric and associated with the “bad” diquarks. This will play a role in the quark models discussed below.
@@ 3.2.2   Spatial wave functions and the non-relativistic harmonic oscillator
To classify the baryon spectrum, we must also discuss the spatial or orbital wave functions $|\psi_{\rm space}\rangle$. The three-body problem is characterized by the time-dependent position of three particles. Instead of the position vectors $\bm{x}_{i}\,(i=1,2,3)$ of the three particles, one usually defines the Jacobi coordinates
@@
which include the trivial center-of-mass (CM) motion and two oscillations represented by the variables $\bm{\rho}$ and $\bm{\lambda}$. The $\bm{\rho}$ oscillator describes the dynamics within a diquark and the $\bm{\lambda}$ oscillator the motion of the third quark relative to the diquark. Permutation symmetry ensures that all three quarks have identical functions. Instead of Jacobian coordinates, some authors Giannini:2001kb , Giannini:2002vp , Giannini:2003xx , DeSanctis:2005kt , Santopinto:2010zz , Santopinto:2012nq , Salehi:2013qga , Giannini:2015zia , Giannini:2016jta , Sattari:2019kkd , Tazimi:2021azh use hypercentral coordinates given by the collective size variable $\sqrt{\rho^{2}+\lambda^{2}}$ and an angle $\arctan\frac{\rho}{\lambda}$.
In a series of papers, Isgur and Karl studied masses and mixing angles of low-mass baryons in a model that contains a flavor-independent confinement potential and hyperfine interactions between quarks due to one-gluon exchange Isgur:1977ef , Isgur:1978xj , Isgur:1978xi , Isgur:1978wd , Isgur:1978xb , Isgur:1979be , Isgur:1979ee . The confining potential is assumed to be harmonic, and the deviation due to the linearly rising confining potential expected in QCD is treated perturbatively. In our following interpretation of the baryon resonance spectrum, we will often use the symmetry properties of the nonrelativistic quark model with a harmonic oscillator as confining potential since it predicts the number of states and, at least qualitatively, their masses. The eigenfunctions of the Hamiltonian are then expanded into eigenfunctions of the harmonic oscillator serving as a complete orthonormal set of basis functions.
@@
Figure 3.2: (color online) Light baryon spectrum in the harmonic oscillator model with band quantum number N, spin $S$, orbital angular momentum $L$, total angular momentum $J$ and parity $P$. $F$ denotes the flavor multiplet (NN for octet, $\Delta$ for decuplet) and $\mathbf{D}$ the spin-flavor multiplet. The assignment with experimental states is discussed in the text. Three- and four-star resonances are shown in bold font. the $J^{P}$ quantum numbers in gray do not have good experimental candidates. The background colors refer to the orbital angular momentum $L$.
The wave functions within a given shell can be further rearranged to form combinations with definite S3 symmetry, as shown in the right panel of Fig. 3.1 (see e.g. Ref. Flamm:1982jv for details). This yields the spatial wave functions $\mathcal{S}_{o}$, $\mathcal{D}_{o}$ and $\mathcal{A}_{o}$ appearing in the left panel of Fig. 3.1. The total orbital angular momentum $\bm{L}=\bm{l}_{\rho}+\bm{l}_{\lambda}$ can adopt values $L=|l_{\rho}-l_{\lambda}|,\dots,l_{\rho}+l_{\lambda}$, and $\bm{L}$ combines with the quark spin $\bm{S}$ to the spin of the particle $\bm{J}=\bm{L}+\bm{S}$. The quantum numbers $J^{P}$ are then obtained from the parity $P=(-1)^{l}$.
@@ 3.3  Light baryon spectrum
Figure 3.4: The number of 3* and 4* $N^{*}$ and $\Delta^{*}$ resonances in ParticleDataGroup:2024cfk above the ground states in 50 MeV bins. The arrows mark the largest spin up to the given mass.
In each cluster there is a state with maximal spin. These states are referred to as stretched states, where the total angular momentum is $J=L+S$. In the high-mass region, the stretched states dominate the reaction, either on physical grounds or because the largest partial wave is the easiest one to identify in angular distributions. The left panels in Fig. 3.5 show the squared baryon mass with maximal spin $J$ in a given cluster, for positive-parity $N^{*}$ and $\Delta^{*}$ resonances, respectively, as a function of $J$. One observes a Regge behavior, i.e., the squared masses of the stretched states depend linearly on $J$ Veneziano:1968yb . Similar Regge trajectories can be drawn for isoscalar mesons, including the $\omega$, $f_{2}(1270)$, $\omega_{3}(1670)$, $f_{4}(1950)$, $\rho_{5}(2350)$, and $f_{6}(2510)$. Mesons and $\Delta^{*}$ baryons have very similar Regge slopes: writing $M^{2}=M_{0}^{2}+aJ$, one finds $a=1.14(2)$ GeV2 for mesons and $a=0.98(2)$ GeV2 for baryons. In mesons, the quark and antiquark are bound by the strong force between a colored quark and an antiquark carrying anti-color. Two colored quarks in a baryon can form a diquark carrying the anti-color of the third quark. The similarity of the meson and baryon trajectories has led to the hypothesis that baryon resonances (more precisely, the stretched states) can be understood as quark-diquark oscillations.
In the following we assign the $N^{*}$ and $\Delta^{*}$ resonances to $SU(6)$ multiplets with defined shell number N, intrinsic orbital angular momentum $L$, total quark spin $S$ and parity $P$. To this end, we combine the masses of baryons in the same multiplet (which differ only by $J$) to the center-of-gravity (c.o.g.) mass [^1] The c.o.g. mass can be used for heavy quarkonia to estimate the mass of the singlet meson from the three triplet states, provided spin-spin and tensor interactions are small. It has been applied to light-quark mesons Klempt:2021nuf even though this is not really justified. The idea is to find a quantity that does not dependent on spin-orbit interactions. defined by